OpenStudy (anonymous):

The area of the square is 25 square meters. Find the measure of the diagonal of the square..

OpenStudy (angela210793):

x^2=25 x=5=side diagonal xsqr2=5sqr2 u can use Pythagorean formula also

OpenStudy (anonymous):

can u show the complete solution and make an explanations....^^

OpenStudy (dumbcow):

Area of square = s^2 -> s^2 = 25 s = sqrt(25) s = 5 The diagonal makes a right triangle with 2 sides of the square Using Pythagorean Theorem a^2 + b^2 = c^2 a =b =5 c^2 = 5^2 + 5^2 c^2 = 25 +25 c^2 = 50 c = sqrt(50) c = sqrt(25)*sqrt(2) c = 5sqrt(2)

OpenStudy (anonymous):

can u show the checking if the answer is correct....ty